# mystyle backgroundcolor=mygray, leftmargin=0.45em, rightmargin=0.45em, innertopmargin=10pt, innerbottommargin=10pt, skipabove=10pt, skipbelow=10pt

### v2

@@ IV-B2 Experimental Results
TABLE IX: Example of Generated Explanation on SeVC.
IV-B3 Case Study
To more intuitively illustrate the importance of evaluating the quality of vulnerability explanations, we analyze an example generated by our model. As shown in Figure IV-B2, this vulnerability code is sourced from SeVC and pertains to pointer-type vulnerabilities, demonstrating incorrect memory deallocation in C language using the delete operator. The explanation adheres to our output paradigm, generating vulnerability detection, location, and detailed information.
The provided explanation correctly identifies the core issue: the data pointer is allocated on the stack using the ALLOCA function but is improperly freed using the delete operator. The explanation points out that the delete operator should only be used to free memory allocated on the heap via the new operator. Using it to free stack-allocated memory can lead to undefined behavior and potential crashes, meeting the accuracy requirement. Additionally, the explanation offers recommendations for addressing this vulnerability, such as using the free function to release stack-allocated memory or using the new operator for heap allocation, and ensuring pointer safety. This satisfies the actionability requirement. The explanation is clear, comprehensive, logically coherent, and easy to read, meeting the clarity requirement.
During the review, we also observe some instances with insufficient quality of explanations. Issues include failure to identify the main vulnerability, vague location, verbose expression, and lack of remediation suggestions. Hallucinations in explanation generation are relatively rare but primarily involved memorized descriptions of similar code samples from the training set and location confusion. Overall, the fine-tuned model demonstrates a certain capability for generating vulnerability explanations.
Summary of RQ2: We find that LLMVulExp can generate explanations of the vulnerability with high Accuracy, Clarity, and Actionability. We also explore the potential of automated data annotation using LLMs to mitigate the manual effort.
@@ IV-C RQ3: How do explanations affect the results of vulnerability detection?
Considering the detection task as a step in the explanation task, we have conducted a comprehensive evaluation of the detection capabilities of the vulnerability explanation model in RQ1. This leads to an important research question: What is the impact of fine-tuning for vulnerability explanation on the detection capabilities of LLMs compared to fine-tuning specialized detection models? In this RQ, we aim to experimentally explore the influence of explanatory information on the performance of vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory information.
Experimental Setup. We conduct our research as ablation studies under the settings of multi-type vulnerability detection scenarios in SeVC (RQ1.3) and multi-label CWE detection scenarios in DiverseVul (RQ1.4). We transform the vulnerability explanation fine-tuning task into a single vulnerability detection task by removing the explanatory information from the training data annotations and modifying the task description and instructions in the prompt. By doing so, we only require the LLMs to output the corresponding vulnerability types and then fine-tune the LLMs to obtain a vulnerability detection model. We compare the detection accuracy of the vulnerability detection model (referred to as Detector) with that of the vulnerability explanation model (referred to as Explainer).
Figure 4: Performance Comparison between Explainer and Detector.
Experimental Results. We present the experimental results for comparison between the Explainer and the Detector on both datasets in Figure 4. The prefix of “Mi” and “Ma” refers to Micro and Macro, respectively. Based on the experimental results, incorporating vulnerability explanation information into the fine-tuning process did not lead to a significant decline in detection performance. For example, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while 74.8% for both on DiverseVul. In fact, it even resulted in performance improvements under certain metrics on DiverseVul (e.g., the Explainer’s Macro-F1 score is 73.2%, which is nearly 2% higher than that of the Detector.).
These observations indicate that the vulnerability explanation task can coexist with the vulnerability detection task without compromising the model’s accuracy in detecting vulnerabilities. The model’s detection capability remains intact with an additional focus on explanation capabilities. Moreover, the automatically annotated vulnerability explanation data contains sufficient domain-specific knowledge, aiding the model in better understanding and identifying various vulnerability patterns. This suggests that enhancing a model’s explanation capabilities can be achieved without sacrificing detection performance. Furthermore, well-annotated explanatory data can improve the model’s overall understanding and performance in both vulnerability detection and explanation tasks.
Summary of RQ3: Integrating vulnerability explanation into the fine-tuning process does not compromise detection capabilities and may even lead to improved performance for certain vulnerability scenarios.
@@ IV-D RQ4: How does the key code extraction impact detection performance?
In this RQ, we aim to investigate if the LLMs can identify the key code that might be prone to vulnerability and further examine the code with a specific focus. To achieve this, we have conducted annotations of key code extraction on both the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the key code information.
Experimental Setup. Similar to the process of annotation discussed in Section III, we use GPT-3.5 to extract the key code. To prevent label leakage, the code vulnerability type tags were not visible during the annotation process. For training data that does not output the original code statements, we will nullify its key code information. If it is test data, it will not be included in the evaluation. During fine-tuning and inference, we modify the prompts used, focusing on task description and instruction sections, to guide the model in focusing on the extracted key code for detecting the vulnerability.
TABLE X: F1 Scores of Key Code Extraction on SeVC (RQ4).
@@
Experimental Results. We present the experimental results of utilizing Key Code Extraction on SeVC in Table X and DiverseVul in Figure 5. We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type). For DiverseVul, there is a back-and-forth trend comparing the results with key code and without key code. The potential reason might be that the longer and more complex code length makes it difficult to effectively extract the key code on DiverseVul. The results indicate that this enhancement scheme can considerably improve the detection accuracy of the model in different vulnerability explanation tasks. This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models. On the one hand, by extracting key portions of the code that are more likely to contain vulnerabilities, the model can concentrate on the most pertinent information, leading to more accurate detection results. On the other hand, key code extraction helps eliminate irrelevant parts of the code, reduce noise, and improve the signal-to-noise ratio. This makes it easier for the model to learn and detect patterns associated with vulnerabilities. This method does not rely on manual feature extraction, making it broadly applicable. Our experimental results preliminarily demonstrate the feasibility of key code extraction through LLMs.
Figure 5: Performances of Utilizing Key Code Extraction on DiverseVul Dataset(RQ4), the numbers in x-axis represent the CWE IDs.
Summary of RQ4: By guiding the LLMs to focus on key code, the performance of vulnerability detection can be improved by a large margin (e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
@@ V Discussion
Figure 6: Performance of Multi-Type vulnerability detection using LLMs with different sizes.
@@ VI Threats to Validity
Construct Validity. We evaluated the explanatory capability using three proposed metrics: Accuracy, Clarity, and Actionability. While these metrics effectively characterize important aspects of vulnerability explanations, they may not comprehensively cover all features. More comprehensive and refined evaluation metrics for vulnerability explanations require further research. To mitigate potential bias, we employed both manual and automated reviews using LLMs. Two authors independently annotated the explanations and resolved any discrepancies until a consensus was reached. The value of Cohen’s Kappa [54] in this process was 0.76, indicating substantial agreement. This rigorous approach helps ensure the reliability and validity of our evaluation.
Internal Validity. Due to the high computational cost of fine-tuning and inference evaluation of LLMs, we were unable to conduct experiments under different data splits and randomization states. The randomness in our experiments (e.g., data splitting, explanation annotation, and fine-tuning process) may affect the results. However, we conducted multiple trials to validate the stability of our conclusions for each experimental setup. Although the number of hyperparameter trials for LoRa configuration and the LLM generation method was relatively limited, and due to GPU constraints, we did not attempt to fine-tune larger models, such as the 34B version, we ensured that these limitations did not fundamentally undermine our primary objective: to explore the potential of conducting vulnerability detection and explanation tasks using LLMs.
External Validity. We chose the advanced open-source code LLM CodeLlama, based on the Llama2 architecture, as our primary model for this research. We tested different versions, including 7B and 13B, and evaluated the performance of the recently released general large language model Llama3 on vulnerability detection and explanation tasks. However, there are many different architectures of general-purpose open-source LLMs and code models, and our experiments were limited to a few models. Additionally, we conducted experiments using two differently configured open-source project-based vulnerability datasets. Although these datasets include a relatively large amount of vulnerable code, there are still differences compared to code in real development environments. Consequently, our experimental conclusions have an inherent risk of not generalizing to other vulnerability datasets or industrial environments. Nonetheless, our careful selection and thorough evaluation process aimed to ensure that our findings remain relevant and insightful within the scope of our study.

### v3

@@ IV-B RQ2: How proficient are LLMs in explaining the detected vulnerabilities?
TABLE VIII: Example of Generated Explanation on SeVC.
IV-B3 Illustrative Example.
To highlight the importance of evaluating vulnerability explanations, we analyze an example from SeVC. The sample pertains to a pointer-type vulnerability caused by incorrect memory deallocation in C code. As shown in Figure IV-B2, the explanation correctly identifies the issue: the delete operator is used to free stack-allocated memory created by the ALLOCA function, which should instead use the free function. The explanation provides actionable suggestions, such as ensuring pointer safety by using the appropriate deallocation methods. It is clear, logically structured, and actionable, satisfying all three evaluation criteria.
Despite these strengths, some generated explanations exhibit issues, such as vague locations, verbose descriptions, or lack of actionable suggestions. While rare, hallucinations may occur, typically involving memorized examples or misinterpreted contexts. Overall, the fine-tuned model demonstrates robust capabilities for generating high-quality vulnerability explanations.
Summary of RQ2: We find that LLMVulExp is capable of generating vulnerability explanations with high levels of Accuracy, Clarity, and Actionability. Additionally, we explore the potential of leveraging LLMs for automated data annotation, significantly reducing the reliance on manual effort.
@@ IV-C RQ3: How do explanations affect the results of vulnerability detection?
Building on the detection capabilities evaluated in RQ1, we consider the detection task as a foundational step in the broader explanation task. This raises an important question: How does fine-tuning for vulnerability explanation influence the detection performance of LLMs compared to models fine-tuned solely for detection? In this RQ, we experimentally examine the impact of incorporating explanatory information on vulnerability detection by comparing the performance of fine-tuned LLMs with and without explanatory data.
Experimental Setup. We conduct ablation studies under the multi-type vulnerability detection scenario in SeVC (RQ1.2) and the multi-label CWE detection scenario in DiverseVul (RQ1.3). To isolate the effect of explanatory information, we transform the vulnerability explanation fine-tuning task into a detection-only task by removing explanatory data from annotations and adjusting task descriptions and prompts. The resulting LLMs output only vulnerability types, creating a detection-focused model (referred to as Detector). This is compared against the performance of the original explanation-focused model (referred to as Explainer).
Figure 5: Performance of Explainer and Detector.
Experimental Results. The comparative results between Explainer and Detector are illustrated in Figure 5, where ”Mi” and ”Ma” denote Micro and Macro metrics, respectively. Incorporating explanatory data into the fine-tuning process did not significantly reduce detection performance. For instance, the weighted precision on SeVC for the Explainer and Detector are 80.1% and 81.1%, respectively, while both achieve 74.8% on DiverseVul. Interestingly, certain metrics improved when explanatory data was included, such as the Macro-F1 score on DiverseVul, where the Explainer outperformed the Detector by nearly 2% (73.2% vs. 71.4%).
These findings demonstrate that the vulnerability explanation task can coexist with the detection task without compromising the model’s detection accuracy. The Explainer retains strong detection performance while gaining additional explanatory capabilities. Furthermore, automatically annotated explanatory data introduces domain-specific knowledge, enhancing the model’s ability to understand and identify diverse vulnerability patterns. In summary, enhancing a model’s explanatory capabilities does not come at the cost of detection performance. In some cases, well-annotated explanatory data can even improve the model’s overall understanding and effectiveness across both tasks.
Summary of RQ3: Integrating vulnerability explanations into the fine-tuning process preserves detection capabilities and may even enhance performance in specific vulnerability scenarios.
@@ IV-D RQ4: How does the key code extraction impact detection performance?
In this RQ, we investigate whether LLMs can identify key code segments prone to vulnerabilities and enhance detection by focusing on these critical sections. To explore this, we annotate key code segments in the SeVC and DiverseVul datasets and fine-tuned the vulnerability explanation LLMs using the annotated key code information.
Experimental Setup. As outlined in Section III, we employ GPT-3.5 to annotate key code segments closely related to vulnerabilities. To avoid label leakage, vulnerability-type tags are concealed during the annotation process. For training samples where key code information was unavailable, the annotations are nullified. Similarly, test samples without key code annotations are excluded from the evaluation. During fine-tuning and inference, the prompts are modified to emphasize task descriptions and instructions, guiding the model to focus on the extracted key code for vulnerability detection.
TABLE IX: F1 of Key Code Extract. on SeVC: (RQ4).
@@
Experimental Results. We present the experimental results of utilizing Key Code Extraction on SeVC in Table IX and DiverseVul in Figure 6. We find that there is a noticeable improvement for SeVC (i.e., 98.5% vs 92.4% for Single-Type and 83.8% vs 64.4% for Multi-Type). For DiverseVul, there is a back-and-forth trend comparing the results with key code and without key code. The potential reason might be that the longer and more complex code length makes it difficult to effectively extract the key code on DiverseVul. The results indicate that this enhancement scheme can considerably improve the detection accuracy of the model in different vulnerability explanation tasks. This demonstrates the importance of supplementing code semantic information for the fine-tuning of large vulnerability models and the feasibility of semantic information extraction based on large models. Key code extraction offers two major benefits: (1) Enhanced Focus: By isolating the most relevant code segments, the model can focus on critical information, leading to more accurate detection results. (2) Noise Reduction: Eliminating irrelevant code improves the signal-to-noise ratio, enabling the model to better learn and detect vulnerability patterns. Moreover, this approach relies on automated annotation rather than manual feature extraction, making it broadly applicable and scalable. Our experimental results preliminarily validate the feasibility and utility of key code extraction through LLMs.
Figure 6: Performance of Key Code on DiverseVul(RQ4).
Summary of RQ4: By guiding LLMs to focus on key code, the performance of vulnerability detection can be significantly improved(e.g., from 64.4% to 83.8% for Multi-Type detection on SeVC).
@@ V Discussion
Figure 7: Performance of different base LLMs.
@@ VI Threats to Validity
Construct Validity. We evaluate explanatory capability using three effective metrics, but these metrics may not cover all aspects of explanations. To mitigate bias, we use both manual and automated reviews with LLMs. Two authors independently annotate the explanations, resolving discrepancies until consensus is reached, achieving a Cohen’s Kappa [46] value of 0.76, indicating substantial agreement.
Internal Validity. Due to the high computational cost of fine-tuning and inference evaluation, we couldn’t conduct experiments under different data splits and randomization states. Although we limit the number of hyperparameter trials for LoRa configuration and didn’t fine-tune larger models like the 34B version due to GPU constraints, these limitations don’t undermine our objective: exploring the potential of LLMs for vulnerability detection and explanation.
External Validity. We choose the advanced open-source code LLM CodeLlama and evaluate different versions as well as the recently released Llama3 on two distinct datasets. These datasets, while containing a significant amount of vulnerable code, include only a few CWE types and mostly generic vulnerabilities, which do not fully reflect the diversity and complexity of vulnerabilities found in real-world development environments. Although these datasets contain substantial vulnerable code, they differ from real development environments, posing a risk of limited generalization.